# Hybrid satellite-fiber quantum network

### v1

@@ Abstract
The length of each link is approximated by the Euclidean distance between the end nodes. The distribution of link lengths shows that $99.84\%$ of all the links are shorter than $61.7$ km [Fig. 3(c)]. This distance corresponds to the intersection in Fig. 2(a) and thus, to achieve optimal distribution rates, intermediate repeaters are added to longer links [marked green in Fig. 3(a)]. The distribution of shortest fiber-length paths connecting randomly selected end nodes Alice and Bob pairs shows that the vast majority of pairs are separated by more than 1 00010001\,0001 000 km [Fig. 3(d)].
@@
The communication between end nodes in the ground network is assumed to take place through the optical fiber path with the shortest length. Along this path, a greedy algorithm is applied to find the optimal placement of both photon and trapped-ion repeaters (i.e., determining which repeaters should be active or bypassed as well as the role to be played by the active repeaters). In this optimal placement, we maximize the end-to-end fidelity or the entanglement distribution rate, depending on the question under consideration, under the constraint of only using repeaters the network is instrumented with (the optimal placement will thus be generally unevenly distributed along the shortest path between the end nodes). Figure 3(a) highlights the shortest fiber path for a choice of Alice and Bob, which has $547.5$ km and $132$ intermediate nodes, where the optimization of the end-to-end fidelity leads to the optimal placement of only 5 photon repeaters (alternated by 4 trapped-ion repeaters).
@@
To devise a hybrid network, we now compare the exclusive use of optical fibers just discussed with the exclusive use of MEO satellites. In the latter case, and only for the purpose of this comparison, each end node would have to be equipped with a ground station (a telescope) able to detect entangled photons. The satellite is then assumed to generate entangled photon pairs and send them directly to the ground stations of Alice and Bob. The photon source on the satellite is considered to generate $N_{\text{s}}=6\times 10^{7}$ entangled down-converted photon pairs per second, where the wavelength is $\lambda=810$ nm and the beam width is $w_{0}=0.15$ m. The radius of the satellite ground station telescopes is taken to be $a_{\text{r}}=0.5$ m (see Supplemental Material [39] for the performance improvement enabled by larger $a_{\text{r}}$). The other parameters are taken as $\eta_{\text{tran}}$ = $\eta_{\text{rec}}$ = $0.5$, $\eta_{\text{filt}}$ = $\eta_{\text{dec}}$ = 0.9, and $\eta_{\text{coup}}$ = 0.7.
Figure 4(a) compares the entanglement distribution rates using only optical fibers or only MEO satellites, for the parameters above in each case. To facilitate comparison, we consider the rate a function of the geographical distance $D$ (rather than fiber length) between the two end nodes. The satellite entanglement distribution rate $R_{\text{s}}$ is almost constant over $D$, at approximately 3 s-1. While using optical fibers, entanglement distillation is performed for a target end-to-end entanglement with a fidelity equal to the satellite fidelity, $F_{\text{s}}$. The distribution rate decreases as the end-to-end geographical distance $D$ increases and falls below the satellite rate when $D>D^{\star}_{\text{R}}\approx 280$ km. On the other hand, Fig. 4(b) shows the fidelity of the end-to-end entanglement via fibers $F_{\text{f}}$ when the time for establishing one copy of entanglement between Alice and Bob is fixed (to be the average time needed to establish one entanglement pair via satellite). The end-to-end fidelity via fibers decreases fast as $D$ increases and drops below $F_{\text{s}}$ for $D>D^{\star}_{\text{F}}\approx 280$ km and below $0.5$ for $D>1200$ km. Figure 4(c) summarizes the relationship between the distribution rate (for a given target fidelity) and the fidelity (for a fixed time) for all Alice-Bob pairs in Fig. 4(a)-(b). The distribution rate increases monotonically with fidelity, as both are strongly correlated with the distance $D$ between the Alice and Bob nodes.
Figure 6: Performance of the hybrid network protocol. Three grid sizes are considered for the placement of satellite ground stations: $d=700$ km (top row), $d=450$ km (middle row), and $d=240$ km (bottom row). (a)-(c) Map of the ground network where the nodes are color-coded according to their distance to the closest satellite ground station (black dots). Distances are marked relative to the point with a latitude ${}^{\text{o}}$ N and a longitude ${}^{\text{o}}$ W. (d)-(f) Scattered plot (left) and histogram (right) of the end-to-end fidelities for the hybrid protocol. (g)-(i) Scattered plot (left) and histogram (right) of the entanglement distribution rates for the hybrid protocol for a target fidelity of 0.87. In (d-i), the repeaters are optimally placed as in Fig. 4. In all histograms, the dashed lines indicate the median.
@@
The results above clearly show that not all pairs of end nodes can communicate most efficiently through optical fibers only or satellites only: the ground network is more efficient at short distances while the satellites are superior at longer distances. However, the distances $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$ at which the relative efficiency inverts depend on the placement of the satellite ground stations. In Fig. 4, we considered the limit case in which every node was equipped with one such station, which offers a lower bound estimation for $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$. In a more realistic scenario, we can equip the network with a grid of stations, as shown in Fig. 5 for a grid spacing $d$ of approximately $700$ km (which limits the maximum distance between an end node and the nearest station to $\approx\!495$ km).
@@
Figure 6 shows a visual representation of these results for $d=700$ km, $450$ km, and $240$ km. To facilitate interpretation, the map is now parameterized by geographic distance (rather than latitude and longitude), and the nodes are color-coded by the length of the shortest optical-fiber path to the closest satellite ground station [Fig. 6(a)-(c)]. For a fair comparison, the time required to establish entanglement via satellite is used as the time available for entanglement distribution and distillation (between endpoints or end points and closest satellite ground stations) via optical fibers. The fidelity decreases slowly as the number of satellite ground stations is reduced [Fig. 6(d)-(f)], with a median of $0.71$ for $d=700$ km; this should be contrasted with the fidelity of 0.87 for the idealized case of $d=0$ km in Fig. 4(b). On the other hand, the median of distribution rate decreases rapidly as $d$ is increased [Fig. 6(g)-(i)]. This is the case because of the exponential loss of photons with the length of the path between end nodes mediated by the ground network, which in turn grows linearly with $d$. In addition, the distribution rate varies substantially more widely as $d$ increases [note the logarithm scale in Fig. 6(g)-(i)].
To examine the impact of varying the spacing of the satellite ground stations, we first note that the distances $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$ are now defined by the intersection between optical fibers only and the combined use of optical fibers and satellite (as in the highlighted path in Fig. 5), which are the inflection points in Fig. 6(d)-(i). Table 1 summarizes the trend in $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$ as $d$ is varied. It follows that even when the entire contiguous U.S. is equipped with only $27$ stations, as for $d=700$ km in Fig. 5, the $D^{\star}_{\text{F}}$ and $D^{\star}_{\text{R}}$ remain approximately $800$ and $640$ km, meaning that the vast majority node pairs would communicate through paths that involve both optical fibers and satellites.
@@
$d$ $d/\sqrt{2}$ $D_{\text{F}}^{\star}$ $D_{\text{R}}^{\star}$
  $700$ km   $495$ km   $800$ km   $640$ km
  $450$ km   $318$ km   $580$ km   $530$ km
  $240$ km   $140$ km   $420$ km   $420$ km
  00 km   00 km   $280$ km   $280$ km
Finally, it is important to consider how the ability to generate and manipulate an increased number of entangled photon pairs, to be enabled by future technologies, would impact our results. Suppose the number of ions at each trapped-ion repeater is increased by a factor $\alpha$ and the photon emission rate $N_{\text{s}}$ at the satellite is increased by a factor $\beta$, relative to the rates considered in Figs. 4 and 6. It follows that the corresponding entanglement distribution rates will also increase by the same factor $\alpha$ or $\beta$, as shown in Fig. 7(a), leading to the same intersections depicted in Fig. 4(a) when $\alpha=\beta$. Accordingly, when $\alpha=\beta$ in the hybrid protocol, the end-to-end fidelity in Fig. 6(d)-(f) remains unchanged while the entanglement distribution rate in Fig. 6(g)-(i) increases by the same factor $\alpha$ [as illustrated in Fig. 7(b)]. The latter is significant because the entanglement distribution rate is otherwise low, with an average of 6.3 s-1 for the idealized case in Fig. 4(a). For $\alpha=\beta=100$, this number increases to the respectable figure of $630$ s-1.
@@
The model for a hybrid quantum network conceived in this paper is capable of establishing entanglement for nationwide quantum communication across the contiguous U.S. Our analysis demonstrates that a hybrid satellite-fiber design offers superior large-scale communication performance compared to either an optical-fiber-based or a satellite-based design alone. In our protocol, optical fibers are invoked for efficient entanglement distribution at short distances, while the satellites are required to maintain a nearly constant distribution rate over long distances.
The hybrid model can be extended in several ways. First, our results are based on a uniform grid distribution of satellite ground stations. A strategic placement of the ground stations that accounts for population density would increase the average satellite-based entanglement distribution rate.
@@
Second, our analysis considered only one satellite. Implementing instead an ensemble of satellites could reduce diffraction and atmospheric extinction and increase the rate of satellite-generated entanglements. Third, while we focused on medium Earth orbit satellites, performance can be further enhanced by considering a multilayer architecture in which satellites at different altitudes are used for communication between ground stations at different distances. Fourth, we considered communication within the contiguous U.S., but similar hybrid network and protocol can be implemented at the global scale. The key step for a global design is to include satellite-to-satellite communication [40], in which entangled photons can be transmitted through two or more satellites before reaching the ground. Altogether, these extensions would strengthen the overall results by increasing the entanglement generation rate, improving fidelity, and expanding the possible number of communicating parties.
@@
The development of a practical large-scale quantum network is currently constrained by technological limitations. At this stage, it is not yet clear what network architecture will prevail. We suggest that by focusing on the co-design of network and protocol, as considered here, one can identify the new technologies to be prioritized and the principles from classical communication networks to be leveraged [41]. Ultimately, this approach is relevant not only for the possible development of a quantum Internet but also for the design of quantum sensing networks for high-precision physics.
VII Ackowledgement
This work was supported by ARO MURI Grant No. W911NF-21-1-0325.
Supplemental Material
Hybrid satellite-fiber quantum network
VIII Comparison with Geostationary satellites
An alternative to MEO satellites would be geostationary (GEO) satellites, and it is therefore instructive to compare the results for the two different systems. GEO satellites orbit Earth approximately 36 0003600036\,00036 000 km above the equator and remain nearly stationary relative to a fixed point on the ground. This geosynchronous positioning, explored in classical telecommunications, simplifies the operation of ground stations, including satellite tracking and pointing. Additionally, each GEO satellite can cover over one-third of the Earth’s surface, enabling communication between distant locations with a single satellite.
We assume that each GEO satellite has an entangled photon generator and emits $N_{\text{s}}=6\times 10^{7}$ pairs of entangled photons per second, as assumed for the MEO satellites in the main text. Using a single satellite for the contiguous portion of the United States, the satellite can be positioned at the longitude that minimizes the average satellite-ground distance. This results in an approximate entanglement distribution rate of only $0.02$ s-1, which is almost independent of the distance between the ground stations [Fig. S1(a)]. For reference, the rate decreases fast over distance and becomes lower for distances above $950$ km if only optical fibers are used with the specifications used in the main text. Here too the fidelity of entanglement distributed by the satellite is constant while the fidelity of entanglement distributed by optical fibers decreases with distance [Fig. S1(b)]. As shown in Fig. S2, the performance of the hybrid protocol combining optical fibers and a GEO satellite is significantly lower than for the corresponding hybrid protocol using MEO satellites. This is due to the higher loss of photons resulting from the higher altitude of these satellites.
@@ VIII Comparison with Geostationary satellites
Figure S1: Performance comparison for entanglement distribution protocols using optical fibers versus GEO satellites. (a) Entanglement distribution rate for optical fibers (blue) and a GEO satellite (orange) with a target fidelity of 0.87. (b) End-to-end entanglement fidelity within the time needed for the satellite to establish one entanglement—the color code is the same as in (a).
@@
Figure S2: Performance of the hybrid protocol integrating optical fibers and a GEO satellite. (a) End-to-end entanglement fidelity within the time needed to establish one entanglement through the satellite. (b) Entanglement distribution rate of the hybrid protocol with a target fidelity of 0.87.
IX Varying the Entanglement swapping fidelity
@@
In the main text, we assumed nearly ideal hardware, where the fidelities of generated trapped-ion photon entanglement and entanglement swapping are $F_{0}$=$F_{\text{swap}}^{\text{p}}$=$F_{\text{swap}}^{\text{i}}=0.99$. However, the current technology of entanglement creation and Bell state measurement renders significantly lower fidelities, especially for photon entanglement swapping. Here, we test the impact of lowering $F_{\text{swap}}^{\text{p}}$. Figure S3 shows the fidelity of end-to-end entanglement established by the hybrid protocol in the main text when $F_{\text{swap}}^{\text{p}}=0.9$ and the other parameters remain unchanged. Compared with $F_{\text{swap}}^{\text{p}}=0.99$ in Fig. 6, the end-to-end entanglement fidelity is necessarily lower for $F_{\text{swap}}^{\text{p}}=0.9$ but the average remains above $0.5$. For the satellite ground stations positioned using the same grid sizes, the median end-to-end fidelity changes from $0.71$, $0.76$, and $0.80$ in Fig. 6 to $0.54$, $0.60$, and $0.69$ in Fig. S3.
@@
Figure S3: End-to-end entanglement fidelity of the hybrid protocol in the main text for $F_{\text{swap}}^{\text{p}}=0.9$. (a-c) Scattered plot (left) and histogram (right) of the end-to-end fidelity using increasingly finer grid sizes for the satellite ground stations (top to bottom). The grid sizes are the same as in Fig. 6(a)-(c).
X Varying the ground station telescope size

### v2

@@ Abstract
The length of each link is approximated by the Euclidean distance between the end nodes. The distribution of link lengths shows that $99.84\%$ of all the links are shorter than $61.7$ km [Fig. 3(c)]. This distance corresponds to the intersection in Fig. 2(a) and thus, to achieve optimal distribution rates, intermediate repeaters are added to longer links [marked green in Fig. 3(a)]. The distribution of shortest fiber-length paths connecting randomly selected Alice-and-Bob end-node pairs shows that the vast majority of pairs are separated by more than 1 0001\,000 km [Fig. 3(d)].
@@
The communication between end nodes in the ground network is assumed to take place through the optical-fiber path with the shortest length. Along this path, a greedy algorithm is applied to find the optimal placement of both photon and trapped-ion repeaters (i.e., determining which repeaters should be active or bypassed as well as the role to be played by the active repeaters). In this optimal placement, we maximize the end-to-end fidelity or the entanglement distribution rate, depending on the question under consideration, under the constraint of only using repeaters the network is instrumented with (the optimal placement will thus generally be unevenly distributed along the shortest path between the end nodes). Figure 3(a) highlights the shortest fiber path for a choice of Alice and Bob, which has $547.5$ km and $132$ intermediate nodes, where the optimization of the end-to-end fidelity leads to the optimal placement of only 5 photon repeaters (alternated by 4 trapped-ion repeaters).
@@
To devise a hybrid network, we now compare the exclusive use of optical fibers just discussed with the exclusive use of MEO satellites. In the latter case, and only for the purpose of this comparison, each end node would have to be equipped with a ground station (a telescope) able to detect entangled photons. The satellite is then assumed to generate entangled photon pairs and send them directly to the ground stations of Alice and Bob. The photon source on the satellite is considered to generate $N_{\text{s}}=6\times 10^{7}$ entangled down-converted photon pairs per second, where the wavelength is $\lambda=810$ nm and the beam width is $w_{0}=0.15$ m. The radius of the satellite ground station telescopes is taken to be $a_{\text{r}}=0.5$ m (see Appendix C for the performance improvement enabled by larger $a_{\text{r}}$). The other parameters are taken as $\eta_{\text{tran}}$ = $\eta_{\text{rec}}$ = $0.5$, $\eta_{\text{filt}}$ = $\eta_{\text{dec}}$ = 0.9, and $\eta_{\text{coup}}$ = 0.7.
Figure 4(a) compares the entanglement distribution rates using only optical fibers or only MEO satellites, for the parameters above in each case. To facilitate comparison, we consider the rate as a function of the geographical distance $D$ (rather than fiber length) between the two end nodes. The satellite entanglement distribution rate $R_{\text{s}}$ is almost constant over $D$, at approximately 3 s-1. While using optical fibers, entanglement distillation is performed for a target end-to-end entanglement with a fidelity equal to the satellite fidelity, $F_{\text{s}}$. The distribution rate decreases as the end-to-end geographical distance $D$ increases and falls below the satellite rate when $D>D^{\star}_{\text{R}}\approx 280$ km. On the other hand, Fig. 4(b) shows the fidelity of the end-to-end entanglement via fibers $F_{\text{f}}$ when the time for establishing one copy of entanglement between Alice and Bob is fixed (to be the average time needed to establish one entanglement pair via satellite). The end-to-end fidelity via fibers decreases rapidly as $D$ increases, and it drops below $F_{\text{s}}$ for $D>D^{\star}_{\text{F}}\approx 280$ km and below $0.5$ for $D>1200$ km. Figure 4(c) summarizes the relationship between the distribution rate (for a given target fidelity) and the fidelity (for a fixed time) for all Alice-Bob pairs in Fig. 4(a)-(b). The distribution rate increases monotonically with fidelity, as both are strongly correlated with the distance $D$ between the Alice and Bob nodes.
[…]
@@
The results above clearly show that not all pairs of end nodes can communicate most efficiently through optical fibers only or satellites only: the ground network is more efficient at short distances while the satellites are superior at longer distances. However, the distances $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$ at which the relative efficiency inverts depend on the placement of the satellite ground stations. In Fig. 4, we considered the limit case in which every node was equipped with one such station, which offers a lower-bound estimation for $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$. In a more realistic scenario, we can equip the network with a grid of stations, as shown in Fig. 5 for a grid spacing $d$ of approximately $700$ km (which limits the maximum distance between an end node and the nearest station to $\approx\!495$ km).
@@
Figure 6 shows a visual representation of these results for $d=700$, $450$, and $240$ km. To facilitate interpretation, the map is now parameterized by geographic distance (rather than latitude and longitude), and the nodes are color-coded by the length of the shortest optical-fiber path to the closest satellite ground station [Fig. 6(a)-(c)]. For a fair comparison, the time required to establish entanglement via satellite is used as the time available for entanglement distribution and distillation (between endpoints or between endpoints and the closest satellite ground stations) via optical fibers. The fidelity decreases slowly as the number of satellite ground stations is reduced [Fig. 6(d)-(f)], with a median of $0.71$ for $d=700$ km; this should be contrasted with the fidelity of 0.87 for the idealized case of $d=0$ km in Fig. 4(b). On the other hand, the median of distribution rate decreases rapidly as $d$ is increased [Fig. 6(g)-(i)]. This is the case because of the exponential loss of photons with the length of the path between end nodes mediated by the ground network, which in turn grows linearly with $d$. In addition, the distribution rate varies substantially more widely as $d$ increases [note the logarithmic scale in Fig. 6(g)-(i)].
To examine the impact of varying the spacing of the satellite ground stations, we first note that the distances $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$ are now defined by the intersection between optical fibers only and the combined use of optical fibers and satellite (as in the highlighted path in Fig. 5), which are the inflection points in Fig. 6(d)-(i). Table 1 summarizes the trends in $D^{\star}_{\text{R}}$ and $D^{\star}_{\text{F}}$ as $d$ is varied. It follows that even when the entire contiguous U.S. is equipped with only $27$ stations, as for $d=700$ km in Fig. 5, $D^{\star}_{\text{F}}$ and $D^{\star}_{\text{R}}$ remain approximately $800$ and $640$ km, meaning that the vast majority of node pairs would communicate through paths that involve both optical fibers and satellites.
@@
$d$ (km) $d/\sqrt{2}$ (km) $D_{\text{F}}^{\star}$ (km) $D_{\text{R}}^{\star}$ (km)
$700$   $495$   $800$   $640$
  $450$   $318$   $580$   $530$
  $240$   $140$   $420$   $420$
  0   0   $280$   $280$
Finally, it is important to consider how the ability to generate and manipulate an increased number of entangled photon pairs, to be enabled by future technologies, would impact our results. Suppose the number of ions at each trapped-ion repeater is increased by a factor $\alpha$ and the photon emission rate $N_{\text{s}}$ at the satellite is increased by a factor $\beta$ relative to the rates considered in Figs. 4 and 6. It follows that the corresponding entanglement distribution rates will also increase by the same factor $\alpha$ or $\beta$, as shown in Fig. 7(a), leading to the same intersections depicted in Fig. 4(a) when $\alpha=\beta$. Accordingly, when $\alpha=\beta$ in the hybrid protocol, the end-to-end fidelity in Fig. 6(d)-(f) remains unchanged while the entanglement distribution rate in Fig. 6(g)-(i) increases by the same factor $\alpha$ [as illustrated in Fig. 7(b)]. The latter is significant because the entanglement distribution rate is otherwise low, with an average of 6.3 s-1 for the idealized case in Fig. 4(a). For $\alpha=\beta=100$, this number increases to the respectable figure of $630$ s-1.
@@
The model for a hybrid quantum network conceived in this article is capable of establishing entanglement for nationwide quantum communication across the contiguous U.S. Our analysis demonstrates that a hybrid satellite-fiber design offers superior large-scale communication performance compared to either an optical-fiber-based or a satellite-based design alone. In our protocol, optical fibers are invoked for efficient entanglement distribution at short distances, while the satellites are required to maintain a nearly constant distribution rate over long distances.
@@
The hybrid model can be extended in several ways. First, our results are based on a uniform grid distribution of satellite ground stations. A strategic placement of the ground stations that accounts for population density would increase the average satellite-based entanglement distribution rate. Second, our analysis considered only one satellite. Implementing instead an ensemble of satellites could reduce diffraction and atmospheric extinction and increase the rate of satellite-generated entanglements. Third, while we focused on medium Earth orbit satellites, performance can be further enhanced by considering a multilayer architecture in which satellites at different altitudes are used for communication between ground stations at different distances. Fourth, we considered communication within the contiguous U.S., but similar hybrid network and protocol can be implemented at the global scale. The key step for a global design is to include satellite-to-satellite communication [38], in which entangled photons can be transmitted through two or more satellites before reaching the ground. Altogether, these extensions would strengthen the overall results by increasing the entanglement generation rate, improving fidelity, and expanding the possible number of communicating parties.
@@
The development of a practical large-scale quantum network is currently constrained by technological limitations. At this stage, it is not yet clear what network architecture will prevail. We suggest that by focusing on the codesign of network and protocol, as considered here, one can identify the technologies that should be prioritized and the principles from classical communication networks that should be leveraged [39]. Ultimately, this approach is relevant not only for the possible development of a quantum internet but also for the design of quantum sensing networks for high-precision physics.
VII Ackowledgements
This work was supported by ARO MURI Grant No. W911NF-21-1-0325. The views and opinions expressed herein are those of the authors and do not necessarily reflect those of the United States Government or any of its agencies.
VIII Data availability
The data that support the findings of this article are openly available [40].
IX Appendix A: Comparison with geostationary satellites
@@
Figure 8: Performance comparison for entanglement distribution protocols using optical fibers versus GEO satellites. (a) Entanglement distribution rate for optical fibers (blue) and a GEO satellite (orange) with a target fidelity of 0.87. (b) End-to-end entanglement fidelity within the time needed for the satellite to establish one entanglement—the color code is the same as in (a).
An alternative to MEO satellites would be GEO satellites, and it is therefore instructive to compare the results for the two different systems. GEO satellites orbit Earth approximately 36 00036\,000 km above the equator and remain nearly stationary relative to a fixed point on the ground. This geosynchronous positioning, explored in classical telecommunications, simplifies the operation of ground stations, including satellite tracking and pointing. Additionally, the high altitude of the GEO satellites results in a large coverage of the Earth’s surface, enabling communication between distant locations with a single satellite.
We assume that each GEO satellite has multiple entangled photon generators and emits $N_{\text{s}}=6\times 10^{7}$ total pairs of entangled photons per second, as assumed for the MEO satellites. Using a single satellite for the contiguous portion of the United States, the satellite can be positioned at the longitude that minimizes the average satellite-ground distance. This results in an approximate entanglement distribution rate of only $0.02$ s-1, which is almost independent of the distance between the ground stations [Fig. 8(a)]. For reference, the rate decreases rapidly over distance and becomes lower for distances above $950$ km if only optical fibers are used with the specifications assumed throughout the article. Here, too, the fidelity of entanglement distributed by the satellite is constant, while the fidelity of entanglement distributed by optical fibers decreases with distance [Fig. 8(b)]. As shown in Fig. 9, the performance of the hybrid protocol combining optical fibers and a GEO satellite is significantly lower than for the corresponding hybrid protocol using MEO satellites. This is due to the higher loss of photons resulting from the higher altitude of these satellites.
@@
Figure 9: Performance of the hybrid protocol integrating optical fibers and a GEO satellite. (a) End-to-end entanglement fidelity within the time needed to establish one entanglement through the satellite. (b) Entanglement distribution rate of the hybrid protocol with a target fidelity of 0.87.
X Appendix B: Varying the Entanglement-swapping fidelity
@@
Throughout the article, we assume nearly ideal hardware, where the fidelities of generated trapped-ion photon entanglement and entanglement swapping are $F_{0}$$F_{\text{swap}}^{\text{p}}$$F_{\text{swap}}^{\text{i}}=0.99$. However, the current technology of entanglement creation and Bell-state measurement renders significantly lower fidelities, especially for photon entanglement swapping. Here, we test the impact of lowering $F_{\text{swap}}^{\text{p}}$. Figure 10 shows the fidelity of end-to-end entanglement established by the hybrid protocol when $F_{\text{swap}}^{\text{p}}=0.9$ and the other parameters remain unchanged. Compared with $F_{\text{swap}}^{\text{p}}=0.99$ in Fig. 6, the end-to-end entanglement fidelity is necessarily lower for $F_{\text{swap}}^{\text{p}}=0.9$ but the average remains above $0.5$. For the satellite ground stations positioned using the same grid sizes, the median end-to-end fidelity changes from $0.71$, $0.76$, and $0.80$ in Fig. 6 to $0.54$, $0.60$, and $0.69$ in Fig. 10.
@@
Figure 10: End-to-end entanglement fidelity of the hybrid protocol in the main text for $F_{\text{swap}}^{\text{p}}=0.9$. (a-c) Scattered plot (left) and histogram (right) of the end-to-end fidelity using increasingly fine grid sizes for the satellite ground stations (top to bottom). The grid sizes are the same as in Fig. 6(a)-(c).
XI Appendix C: Varying the ground station telescope size